OpenStudy (anonymous):

Graph the linear equation by solving For y first x+5y=15

OpenStudy (anonymous):

To solve for y simply isolate it So you want to move x to the other side 5y = 15 - x Divide 5 y = (15-x)/3

OpenStudy (anonymous):

y=3-1/5x

OpenStudy (anonymous):

Sorry yeah, 5 is on the bottom it was a typo. (15-x)/5

OpenStudy (anonymous):

how would I graph it